OpenStudy (anonymous):

9. Lisa can ride her bike to work in 2/3 of an hour. When she drives to work it takes 1/3 of an hour. If she drives 12 miles per hour faster than she rides her bike, how fast does she ride her bike? 12 mph 6 mph 15 mph 18 mph

OpenStudy (anonymous):

let \[v_c=v+12\] \[v_b=v\]

OpenStudy (anonymous):

let v be her "normal" velocity and s be the distance to work s = v * (2/3) s = (v+12) * (1/3) divide the equations and get (v+12) = 2v v = 12

OpenStudy (anonymous):

if \[V=\frac{d}{t}\] \[v*t=d\]

OpenStudy (anonymous):

thanksssss

OpenStudy (anonymous):

since distances are the same \[v*\frac{2}{3}=(v+12)\frac{1}{3}\] \[\frac{2}{3}v=\frac{1}{3}v+4\] \[\frac{1}{3}v=4\] \[v=12\]